OpenStudy (anonymous):

The ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6. I think this is false (ned correcter.... Thanks:--) )

OpenStudy (across):

All you have to do is plug in the numbers and check for yourself: \[1+2(-5)=-9\]\[1-10=-9\]\[-9=-9\] \[1-(-5)=6\]\[1+5=6\]\[6=6\] Both are true, so yes.